A worker at the top of a 526-m-tall television transmitting tower accidentally drops a heavy tool. If air resistance is negligible, how fast (in m/s) is the tool going just before it hits the ground?
1
Expert's answer
2019-04-09T09:02:04-0400

We can find the final velocity of the tool just before it hits the ground from the kinematic equation:


"v_f^2 = v_0^2 + 2gh,"

here, "v_0 = 0" is the initial velocity of the tool, "v_f" is the final velocity of the tool just before it hits the ground, "g" is the acceleration due to gravity, "h" is the height of the television transmitting tower.

Then, we get:


"v_f = \sqrt{2gh} = \sqrt{2 \cdot 9.8 \dfrac{m}{s^2} \cdot 526 m} = 101.5 \dfrac{m}{s}."

Answer:

"v_f = 101.5 \dfrac{m}{s}."